OpenStudy (anonymous):

The equation of a curve is given by y = x^2 + ax + 3, where a is a constant. Given that this equation can also be written as y = (x + 4)^2 + b, find (i) the value of a and b, (ii) the coordinates of the turning point of the curve Checking answer ;; test tomorrow..........

OpenStudy (anonymous):

if this can be written as \[(x+4)^2+b\] then you have \[x^2+8x+16+b=x^2+ax+3\]

OpenStudy (anonymous):

this tells you that \[8=a\] and \[16+b=3\] so \[b=-13\]

OpenStudy (anonymous):

"turning point" (yikes) is the vertex. the graph is a parabola that opens up, so it "turns" at the vertex, when it goes from heading down (decreasing) to heading up (increasing)

OpenStudy (anonymous):

we know that \[b=-13\] and so we have \[y=(x+4)^2-13\] making the vertex (-4,-13) that is, since the first term \[(x+4)^2\] is always greater than or equal to zero, it is smallest when it is 0. it is 0 when x = -4, and if x = -4, then y = -13

OpenStudy (anonymous):

thank you :'D